The spectroscopic data in the table is generated with five solutions of known concentration.

Concentration (M)   Absorbance
0.0183    0.1375
0.0366    0.1650
0.0732    0.5256
0.146    1.038
0.293    2.306
What is the slope of the linear regression line formed by these points?
?=

M−1
What is the intercept of the linear regression line?
?=

What is the concentration of a solution with an absorbance of 0.383 ?
concentration =

M
What is the concentration of a solution with an absorbance of 1.109 ?
concentration =

M
Answer #1

Answer:

The plot concentration vs absorbance is

Form the above plot, the straight line equation is

Absorbance=(8.04046) concentration + (-0.07753) with R^2=0.995.

y=mx + b

Where m=slope and b=y-intercept.

(a) The slope of the line, m=8.04046 M^-1.

(b) The y-intercept is b=-0.07753.

(c) Given absorbance=0.383, then concentration=?

Concentration=(absorbance + 0.07753)/(8.04046 M^-1)

Concentration=(0.383+0.07753)/(8.04046)=0.0572 M.

(d) Given absorbance=1.109, then

Concentration=(1.109+0.07753)/(8.04046)=0.1476 M.

Please let me know if you have any doubt. Thanks and I hope you like it.